A vessel of volume 0.2m3 contains nitrogen at 1.013 bar and 15°C. If 0.2 kg of nitrogen is now pumped into the vessel, calculate the new pressure when the vessel has returned to its initial temperature. The molar mass of nitrogen is 28 kg/K - mol and it may be assumed to be a perfect gas.



1
Expert's answer
2021-07-07T08:38:57-0400

Solution.

"V=0.2m^3;"

"p_1=1.013\sdot10^5Pa;"

"T=288K;"

"m_2=0.2kg;"

"M=28\sdot10^{-3}kg/mol;"

"p_2V=\dfrac{m_2}{M}RT\implies p_2=\dfrac{m_2RT}{MV};"

"p_2=\dfrac{0.2\sdot8.31\sdot288}{28\sdot10^{-3}\sdot0.2}=0.855\sdot10^5Pa;"

"p=p_1+p_2;"

"p=1.013\sdot10^5Pa+0.855\sdot10^5Pa=1.868\sdot10^5Pa;"

Answer: "p=1.868" bar.